Given š‘„ = 60; and š‘  = 6, find the z-score that corresponds to each of the following scores up to two decimal places


1
Expert's answer
2021-03-30T07:49:49-0400


a.


"z=\dfrac{x-\bar{x}}{s}=\dfrac{70-60}{6}=\dfrac{5}{3}\approx1.67"

b.


"z=\dfrac{x-\bar{x}}{s}=\dfrac{58-60}{6}=-\dfrac{2}{6}\approx-0.33"